OpenStudy (anonymous):

If the yearbook staff sells 50 yearbooks, the cost per yearbook is $45. If they sell 125 yearbooks, the cost per yearbook is $37.50. Assuming a linear equation fits the situation, find the price per yearbook if the staff sells 150 yearbooks

OpenStudy (anonymous):

Is total cost linear, or is average cost linear?

OpenStudy (anonymous):

Let fixed cost be F and variable cost be V. Then, cost function is C(x) = Vx + F For 50 books, cost per book is $45. Therefore, 45*50 = V*50 + F For 125 books, cost per book is $37.50. Therefore, 37.50*125 = V*125 + F On solving the above two equations we get V = 32.5 and F = 625 Therefore, cost function is C(x) = 32.5x + 625 When number of books is 150, cost is C(150) = 32.5*150 + 625 = $5500 Cost per book is 5500/150 = $36.67 Hope it helps.

OpenStudy (anonymous):

So its 36 dollars. :)